OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points.P(6, 2), Q(8, -4)

OpenStudy (anonymous):

it has to be in _x_y______ form

OpenStudy (anonymous):

Gradient = (2 - (-4)) / (6 - 8) = -3 Using P(6 , 2): y = 2 - 3(x - 6) = 2 - 3x + 18 = 20 - 3x 3x + y = 20 (answer, I assume this is the form you are trying to get)

OpenStudy (anonymous):

yes, thank you!